OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

Number of human b, number of horse -a a+b=74 4a+2b=196

OpenStudy (anonymous):

Number of human b, number of horse -a a+b=74 4a+2b= 196

OpenStudy (anonymous):

24 horse 50 human

OpenStudy (anonymous):

Wait so which one is right?

OpenStudy (anonymous):

50 human have 100 legs, 24 horse has 96 legs , combined total of 196, Checks out

OpenStudy (anonymous):

Oh sweet. Gracias Amigos